OpenStudy (candy13106):

find the surface area of a rectangular prism that has a length of 8 inches,a width of 3 inches,and a height of 6 inches

OpenStudy (anonymous):

Surface Area of a Rectangular Prism = 2ab + 2bc + 2ac http://www.math.com/tables/geometry/surfareas.htm Can you figure it out? :o)

OpenStudy (candy13106):

noooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooo

OpenStudy (anonymous):

Look at the side with the picture and you'll know which ones are a, b, and c. Then, plug them in and you'll have your answer. :o)

OpenStudy (candy13106):

there is no picture there is just a problem

OpenStudy (candy13106):

i dont get that \

OpenStudy (anonymous):

A rectangular prism has 2 ends and 4 sides. Opposite sides have the same area. The surface area is the sum of the areas of all six sides. Find the area of two sides (Length*Height)*2 sides Find the area of adjacent sides (Width*Height)*2 sides Find the area of ends (Length*Width)*2 ends Add the three areas together to find the surface area Example: The surface area of a rectangular prism 5 cm long, 3 cm. wide and 2 cm. high = 5*2*2 + 3*2*2 + 5*3*2 = 20 + 12 + 30 = 62 cm2.

OpenStudy (anonymous):



OpenStudy (anonymous):

a=3, b=6, c=8 Plug and solve

OpenStudy (candy13106):

this is exactly what i saw on the internet and idk

OpenStudy (anonymous):

Multiply the length and height and then multiply it by 2 so (6*8) * 2

OpenStudy (candy13106):

i did that and the teacher said it was so, wrong

OpenStudy (anonymous):

Then multiply separately the width and height so (3*8) *2

OpenStudy (anonymous):

opps make that a 6 not 8

OpenStudy (anonymous):

Then multiply length and width so (8*3) *2

OpenStudy (candy13106):

all rith thank u so, much

OpenStudy (anonymous):

Then add them together, you have to get the right answer. You're teacher is a dumbass if she says it's wrong.

OpenStudy (anonymous):

a=3, b=6, c=8 2ab+2bc+2ac 2(3)(6)+2(6)(8)+2(3)(8) 36+96+48=180

OpenStudy (candy13106):

i get it i get it